Update for the Week Ending May 27, 2016
Site Remediation Continues
- Continued the application of water and calcium chloride for dust control;
- Continued the application of odor-suppressant foam and BioSolve for odor and dust control;
- Maintained erosion and sediment (E&S) controls (Installed waddles - wood chip and straw filled fabric tubes - around the Site perimeter to prevent silt and sediment from leaving the Site and covered all stockpiles);
- Continued operation of the water treatment plant (WTP) and discharge to the sewer;
- Continued lining and loading trucks and shipping impacted soil to Bayshore Soil Management for treatment and disposal:
- 31 trucks; 865.25 tons shipped last week; and
- 4,024.85 tons shipped to date.
- Continued excavation and installation of the cement/bentonite (CB) slurry trench for installation of the Support of Excavation (SOE) wall sheet piles:
- 126 linear feet installed last week; and
- 282 of 629 linear feet installed to date
- Continued the installation of the sheet piles in the CB slurry trench:
- 96.5 linear feet installed last week; and
- 137.8 of 629 linear feet installed to date.
Recently completed activities:
- Completed installation of the CB slurry trench along the south alignment of the SOE wall.
- Relocate and expand the temporary containment building (TCB) over the SOE wall area in preparation for conducting the deep excavation;
- Install deep wells for dewatering in preparation for conducting the deep excavation; and
- Initiate the deep excavation within the SOE wall.
Community Air Monitoring Plan Summary
Community air monitoring, initiated on Tuesday, December 1, 2015, continues as intrusive (ground disturbance) activities are conducted.
There were no fifteen-minute Action Limit exceedances for Total Volatile Organic Compounds (TVOCs) or odors during this reporting period.
There were two instances of the fifteen-minute average concentration of Particulate Matter (PM-10) [dust], exceeding the applicable Action Limit on May 24, 2016, associated with offloading cement and mixing operations at the grout plant. A plastic shroud and additional water spraying were deployed, and dust levels returned to below the Response Limit.
There was one instance of the fifteen-minute average concentration of Particulate Matter (PM-10) [dust] exceeding the applicable Action Limit on May 20, 2016, associated with the transfer and loading of bentonite at the grout plant. Work was stopped, and the bentonite bags were removed by laborers instead of the fork lift, and dust levels returned to below the Response Limit.
If you have any questions or comments about the work on the Site, please use the Williamsburg MGP Telephone Hotline at (718) 403-3053.
Photos of Activity
Installing sheet piles in the CB slurry trench along the east alignment of the SOE wall.
Excavating deep obstruction during installation of the CB slurry trench along the east alignment of the SOE wall and applying odor control foam.
Excavating soil during installation of the CB slurry trench along the south alignment of the SOE wall.
Excavating and installing the CB slurry trench along the south alignment of the SOE wall.
Offloading excavated CB spoils in the lined containment area in the TCB.
Repairing tears to the fabric caused by wind on the western gable end of the TCB.
Applying foam to CB slurry trench spoils for odor and vapor control in the TCB.